Tuesday 12/03/13

9:15a – 11:15p


  • Call Dad, happy birthday
  • Analyze Ph polymer simulations for meeting with Ting.


  • Finish D07 on STORM2 (9 kb green, transcribed, RNaseA treated). Looks transcribed – variable spots, larger than expected for only 9 kb maybe?
  • Start E02 on STORM2. (13 kb green (between R/Y)) These look better, pretty clear single spots.
  • Scores of beads drifting through the field of view, none stuck. Damn it. I don’t have time to fix this this morning.
  • Replaced beads, Adhered with EDAC. Now running E02 imaging O/N.
  • Hot washes of RNA stains: D12, E08 and E10 RNA.
  • STORM – no signal in D12 RNA. Not good, this worked last time. Maybe RNase contamination?
  • no signal in E08 RNA either
  • Attempting to image E10dna (previously imaged RNA). Can’t switch scopes easily — Steve-array offsets much easier to same system!
  • Broke coverslip removing from STORM 4. Scored coverslips are more delicate, new tape was very strong and the objective was still in contact with the slide.

Meeting with Ting Wu

  • not that excited about HiC comparisons
  • very excited about showing flanking regions
  • try secondary labeling techniques. Design primers for this tomorrow?
  • need some genetic manipulation test. Maybe heat shock, maybe LPS. Or siRNA?

Analyzing dot data

  • list2img not working given xc, yc, zc, and a data. No idea why. Suspect an imaxes problem. Now troubleshooting.
  • Fixed list2img (bug with Z and bug with exclusive inequalities). Still need to test that it works correctly with with STORMrender still.
This entry was posted in Summaries and tagged , . Bookmark the permalink.